OpenStudy (anonymous):

Solve, 1/4 (20 - 4 a) = 6- a 0 no solution -1 all real numbers PLEASE HELP!! D:

OpenStudy (anonymous):

Fractions are hard for me.. :/

OpenStudy (photon336):

this is similar to the distributive property c(a+b) = c*a + c*b \[1/4(20-4a) = (1/4)*(20) - (1/4)*(4a) = (5-a) = 6 -a \]

OpenStudy (anonymous):

no solution

OpenStudy (photon336):

we just distribute the 1/4 then after this we solve \[\frac{ 20 }{ 4 }- \frac{ 4 }{ 4 }a = (5-a) \] \[5-a = 6-a \]

OpenStudy (anonymous):

Kitty449, quick off the topic question.. your profile pic.. is that really you? O_O

OpenStudy (anonymous):

im curious :D

OpenStudy (anonymous):

yes why?

OpenStudy (anonymous):

Smart and pretty.. you got a good life ahead of you.. lol

OpenStudy (anonymous):

thank you lol

OpenStudy (photon336):

there's no solution for this

OpenStudy (anonymous):

thanks!

OpenStudy (photon336):

\[5-a = 6 -a \] 5-6 = -1 = 0 ? nope